A galvanic cellis applied by the combination of fe rodimmersed in feso4 solution of concentration 0.25M


1
Expert's answer
2022-01-31T20:28:01-0500

"Fe^{2+}_{(aq)}+2e^-\to\>Fe_{(s)}" "E^{\theta}=-0.44"

If the cell has Fe rod in FeSO"_4" at concentration of 0.25M and Cu rod in copper solution of 0.45M


Then E.M.F can be calculated as follows;


"E^{\theta}_{cell}=E^0_{Cu^{2+}/Cu}-E^0_{Fe^{2+}/Fe}=0.34-(-0.44)"


"=0.78"



"E^{\theta}_{cell}=0.78+\frac{0.
0592}{2}\>log\>\frac{0.25}{0.45}"


"=0.77V"